# Slab tilings, flips and the triple twist

### v1

@@ Abstract
A domino is a $2\times 1\times 1$ parallelepiped formed by the union of two unit cubes and a slab is a $2\times 2\times 1$ parallelepiped formed by the union of four unit cubes. We are interested in tiling regions formed by the finite union of unit cubes. We investigate domino tilings, slab tilings and mixed tilings (vertical dominoes and horizontal slabs). A flip is a local move: in a slab (resp. domino) tiling, two neighboring parallel slabs (resp. dominoes) are removed and placed back in a different position.
We may interpret a three dimensional region as being made of floors. A slab or domino which crosses two floors is vertical, and horizontal otherwise. We can also define a flip in a mixed tiling, we replace four vertical dominoes in a $2\times 2\times 2$ box by two horizontal slabs and vice-versa.
Inspired by the twist for domino tilings we construct a flip invariant for mixed tilings. Based on this, we then construct the triple twist for slab tilings: an invariant under flips assuming values in ${\mathbb{Z}}^{3}$. We show that if the region is a large box then the triple twist assumes a large number of possible values, at least proportional to the fourth power of the volume. We also give examples of smaller regions for which the triple twist assumes only one value and the set of tilings is connected under flips.
@@ 1 Introduction
We choose to depict a three-dimensional domino tiling by floors, numbered from left to right. The $n$-th floor from left to right corresponds to $z\in[n,n+1]$, $n\in{\mathbb{Z}}$. A domino which is contained in only one floor (named horizontal domino) is represented as a $2\times 1$ light grey rectangle. A domino which crosses two floors (named vertical) is represented as two identical $1\times 1$ squares which have the same projection on the first floor, with the square on the lower floor (left) being painted dark grey and the square on the upper floor (right) being painted white. For convenience, we may at times speak of a region using vocabulary of the floor diagram for the region. That is, we talk of the corresponding square of the floor diagram when referring to a certain cube.
@@
Figure 1: Tiling of a $3\times 3\times 2$ box; on the left the floor $z\in[0,1]$, on the right the floor $z\in[1,2]$. The $x$ and $y$ axes are usually drawn as in this example. In this case $\operatorname{Tw}=-1$.
@@
In order to facilitate visualization we choose to depict a three-dimensional slab tiling by floors, numbered from left to right. A slab which is contained in only one floor (named horizontal) is represented as a $2\times 2$ light grey square and a slab which crosses two floors (named vertical) is represented as two identical planar dominoes which have the same projection on the first floor, with the domino on the lower (left) floor painted dark grey and the domino on the upper floor (right) painted white.
Figure 2: A tiling of a $6\times 6\times 6$ box.
@@
We can define flips for slabs in the same way as for dominoes: we take two slabs which have a $2\times 2$ face in common, remove them from the region, and then place them back in one of the two other positions.
The other objects of interest are mixed tilings. Given a three dimensional region ${\cal R}$, a mixed tiling is a decomposition of ${\cal R}$ into horizontal slabs and vertical dominoes. Note we admit as mixed tilings the configurations where there are only horizontal slabs or only vertical dominoes.
Figure 3: A non trivial example of mixed tiling.
One may notice the mixed tiling in Figure 3 can be obtained from the slab tiling in Figure 2 by slicing each vertical slab in half along a plane perpendicular to the paper. This process of slicing vertical slabs in half to transform a slab tiling into a mixed tiling can always be done. However, it is not always possible to undo this process in a well defined manner. For example, in Figure 3 there are many ways to join the vertical dominoes which cross the first and second floor. Also, there are mixed tilings which do not come from any slab tiling, as in Figure 4.
Figure 4: A mixed tiling which does not correspond to a slab tiling. Obtained from the tiling in Figure 3 by taking the central $4\times 4\times 6$ box.
We can also define a flip in a mixed tiling. In one way, we take two adjacent slabs which lie on different floors, perform a slab flip and then slice the resulting vertical slabs in half. In the other way, we take four vertical dominoes which form a $2\times 2\times 2$ cube, join two pairs of adjacent dominoes and perform a slab flip to end with two horizontal slabs. See Figure 5 for an example.
Figure 5: We show one flip performed on the tiling of Figure 3.
Given two slab tilings of the same three dimensional region ${\cal R}$, we may ask whether it is possible to transform one of them into the other by performing a sequence of flips. There is a straightforward interpretation of this situation in a new graph: vertices are tilings and edges join pairs of tilings which differ by a flip. Depending on the region ${\cal R}$, this graph may split into many connected components, which can sometimes be distinguished by the triple twist. In some simple cases, such as a $4\times 4\times N$ or an $M\times N\times 2$ box, there is only one connected component.
In Section 2 we define and review basic properties of the twist, a flip invariant for domino tilings. In Section 3 we proceed to construct invariants for mixed and slab tilings. In Section 4 we state some properties of these invariants. In Section 5 we study the connected components via flips of the space of tilings. As mentioned above, in some special cases that space is connected; in general the number of connected components is large. Finally, in Section 6 we list some open problems.
The second author is thankful for the generous support of CNPq, CAPES and FAPERJ (Brazil). The first and third authors are thankful to CNPq for the support in the PIBIC undergraduate project.
2 Twist of domino tilings
We start this section by defining the twist for a domino tiling ${\mathbf{t}}$ of a region ${\cal R}$. There are several equivalent definitions; we present an explicit construction, and not in the greatest possible generality. For a more general discussion see [3, 1].
@@ 2 Twist of domino tilings
For that, we first define the effect, a function which takes two dominoes in ${\mathbf{t}}$ and returns a value in $\{0,\pm 1\}$. On ${\mathbf{t}}$, we choose a horizontal domino $d_{0}$ whose centers of the cubes it covers form a line parallel to $e_{2}$. Let $d_{1}$ be a vertical domino in ${\mathbf{t}}$. We say $d_{1}$ affects $d_{0}$ if any of the cubes of $d_{1}$ is on the same floor as $d_{0}$ and if the projections of the interiors of $d_{0}$ and $d_{1}$ over the plane $x=0$ have non-empty intersection. Figure 6 shows some examples for which there are effects. If these conditions do not hold we say the effect of $d_{1}$ on $d_{0}$ is 00.
@@
Figure 6: Examples of effects for pairs of dominoes.
We then define the twist $\operatorname{Tw}({\mathbf{t}})$ as $\frac{1}{4}$ times the sum of all the effects of $d_{1}$ on $d_{0}$ for all the suitable pairs $(d_{0},d_{1})$. For example, the twist of the tiling in Figure 1 is $-1$, since each floor has a contribution of $-\frac{1}{2}$.
We list, without proof, properties of the twist that interest us in this paper. The first property is that it is an invariant under domino flips, which can be checked by casework. The second property is that the twist is integer valued for tilings of cylinders [3] justifying the constant of $\frac{1}{4}$.
The third property is that if ${\cal R}$ is a cylinder, the twist is left unchanged after a rotation in ${\mathbb{R}}^{3}$. The fourth property is that the twist changes sign under reflections by any of the planes $x=0$, $y=0$ or $z=0$, since any of them does not change the quantity of effects and changes exactly one parameter in all effects. Note the resulting floor diagram does not represent a cylinder in general, but does so if ${\cal R}$ is a box. In Figure 7 we show two examples of tilings of cylinders, and in Figure 8 we rotate each of them, noting the twist is unchanged.
Figure 7: Two examples of tilings of cylinders. In both cases, $\operatorname{Tw}=0$.
Figure 8: The $x$-view of the two tilings in Figure 7, i.e., the same tilings after a rotation. Notice that we still have $\operatorname{Tw}=0$.
@@ 3 Coloring and twist
Figure 9: A slab-type coloring of a $6\times 6\times 6$ box, based on Example 3.1.
@@
Figure 10: The top left corner of the first two floors of ${\cal R}$.
@@
The second step of the process is to transform the region. We first consider cyllinders of the form ${\cal R}={\cal D}\times[0,N]$ and the $z$-view. A pair of colors $(c_{0},c_{1})$ is good (for this view) if $c_{0}$ and $c_{1}$ are on a same diagonal on the first floor: the good pairs are therefore red-green and blue-yellow. The colors in a good pair are also called good; the other colors are bad. Informally, we keep all cubes (or squares) of one of the two good colors and delete all cubes of bad colors. We then inflate the remaining cubes so as to obtain cuboids which fill in the space left by the deleted cubes. Figure 11 shows an example: the initial region is as in Figure 9 and the good pair is red-green.
Figure 11: The region in Figure 9 is transformed by deleting the blue and yellow cubes and inflating the red and green ones.
@@
We are now left with the task of transforming the tilings. We first consider slab tilings; let ${\mathbf{t}}$ be a tiling of ${\cal R}$. Each slab $P$ of ${\mathbf{t}}$ covers two preserved cubes and two deleted cubes. Let $S_{0}$ and $S_{1}$ be the two preserved cubes, which are transformed into $\tilde{S}_{0}$ and $\tilde{S}_{1}$, respectively. We say that $P$ is taken to the domino $\tilde{P}$, the union of $\tilde{S}_{0}$ and $\tilde{S}_{1}$. See Figures 2 and 12 for an example.
Figure 12: The domino tiling obtained from the slab tiling in Figure 2 or the mixed tiling in Figure 3. In the second row we show the same tiling rotated clockwise by $45^{\circ}$. We have $\operatorname{Tw}=+2$.
@@
Figure 13: The figures corresponding to the $y$-view, with the blue-green good pair. In this case $\operatorname{Tw}(\tilde{\mathbf{t}})=0$.
@@
Figure 14: The figures corresponding to the $x$-view, with the green-yellow good pair. In this case $\operatorname{Tw}(\tilde{\mathbf{t}})=0$.
@@
Figure 15: Two possibilities of flipping two slabs on the top left corner of the first floor of Figure 2. Under the transformation, the first possibility yields the same domino tiling as in Figure 12.
Figure 16: The domino tiling from the second possibility in Figure 15 after the transformation.
@@ 4 Properties of the Triple Twist
Figure 17: An example of a mixed tiling. We assume the top left cube on the first floor is $[0,1]^{3}$, which is green.
@@
Figure 19: We have, in order: one instance of case 1, two instances of case 2 and two instances of case 3.
@@
The following proposition is similar, but for slab tilings. A symmetry of ${\mathbb{R}}^{3}$ which preserves unit cubes is a map of the form $q(w)=v_{0}+Qw$ where $v_{0}\in{\mathbb{Z}}^{3}$ and $Q$ is a $3\times 3$ permutation matrix with signs. Let $\operatorname{sign}(q)=\det(Q)\in\{\pm 1\}$.
@@
5 On connected components
In this section we study connected components under flips of the space of slab tilings of several kinds of boxes. We begin with two-floored regions.

### v2

@@ Abstract
A domino is a $2\times 1\times 1$ parallelepiped formed by the union of two unit cubes and a slab is a $2\times 2\times 1$ parallelepiped formed by the union of four unit cubes. We are interested in tiling regions formed by the finite union of unit cubes. Domino tilings have been studied before; here we investigate slab tilings. As for domino tilings, a flip in a slab tiling is a local move: two neighboring parallel slabs are removed and placed back in a different position.
Inspired by the twist for domino tilings, we construct a flip invariant for slab tilings: the triple twist, assuming values in ${\mathbb{Z}}^{3}$. We show that if the region is a large box then the triple twist assumes a large number of possible values, roughly proportional to the fourth power of the volume. We also give examples of smaller regions for which the set of tilings is connected under flips, so that the triple twist assumes only one value.
@@ 1 Introduction
We give a brief summary of the directly relevant results about 2d and 3d domino tilings. We choose to depict a three-dimensional domino tiling by floors, numbered from left to right. The $n$-th floor from left to right corresponds to $z\in[n,n+1]$, $n\in{\mathbb{Z}}$. A domino which is contained in only one floor (named horizontal domino) is represented as a $2\times 1$ light grey rectangle. A domino which crosses two floors (named vertical) is represented as two $1\times 1$ squares which have the same projection on the first floor, with the square on the lower floor (left) being painted dark grey and the square on the upper floor (right) being painted white. Figure 1 shows a domino tiling of the $3\times 3\times 2$ box. On the left, the floor $z\in[0,1]$; on the right, the floor $z\in[1,2]$. The $x$ and $y$ axes are usually drawn as in this example. For convenience, we may at times speak of a region using vocabulary of the floor diagram for the region. That is, we talk of the corresponding square of the floor diagram when referring to a certain cube.
@@
Figure 1: Domino tiling of a $3\times 3\times 2$ box
We are interested in local moves. The most natural local move is a flip: remove two parallel dominoes and place them back in the only other possible position. We provide a graph theoretical formulation of this scenario. Given a region ${\cal D}$, construct a new graph, where vertices are domino tilings ${\mathbf{t}}$ of ${\cal D}$ and edges join pairs ${\mathbf{t}}_{0},{\mathbf{t}}_{1}$ of tilings which differ by a flip. If ${\cal D}\subset{\mathbb{R}}^{2}$ is a quadriculated disk, i.e., a bounded 2d contractible region, then this graph is connected. Equivalently, any two tilings of ${\cal D}$ can be joined by a finite sequence of flips [8]. Applied to other regions (possibly 3d), this construction allows us to speak of connected components under flips.
For 3d domino tilings, this new graph is almost never connected. As a first example, consider the box in Figure 1 and the graph whose vertices are its tilings. This box admits $229$ tilings (see [6], which includes larger examples). The tiling in Figure 1 admits no flips (it corresponds to an isolated vertex in the graph). The mirror image is likewise isolated. The other $227$ tilings form a single connected component under flips. It might be conjectured after this example that almost all tilings of a 3d box can be joined by a finite number of flips: as we shall see, this is also not correct.
Given a contractible cubiculated region ${\cal R}\subset{\mathbb{R}}^{3}$, there exists a map $\operatorname{Tw}$ (twist) from the set of domino tilings of ${\cal R}$ to ${\mathbb{Z}}$. There are several ways to construct the function $\operatorname{Tw}$: see [1, 3] for more general constructions. For completeness, we include the construction from [4] in Section 2: that construction is valid if ${\cal R}$ is a cylinder. This map has several important properties. The main property is that twist is an invariant under domino flips, so that connected components under flips are contained in level sets of $\operatorname{Tw}$. For the $3\times 3\times 2$ box, the tiling in Figure 1 has $\operatorname{Tw}=-1$; its mirror image has $\operatorname{Tw}=+1$. The other $227$ tilings all have $\operatorname{Tw}=0$.
Twist also works well with symmetries of the region ${\cal R}$. For instance, with our construction, if ${\cal R}$ is a cylinder then the twist is left unchanged after a rotation in ${\mathbb{R}}^{3}$. Also for cylinders, the twist changes sign under reflections by any of the planes $x=0$, $y=0$ or $z=0$.
For domino tilings, there is another important local move besides the flip: the trit ([4, 1]). In order to perform a trit in a tiling ${\mathbf{t}}$, search for three dominoes whose union is a $2\times 2\times 2$ cube minus two unit cubes in opposite vertices. The trit is then performed by removing the three dominoes and placing them back in the only other possible position. A trit always changes the twist by adding ±$\pm 1$, where the sign depends on orientation. The tiling in Figure 1 admits four possible trits, all of them taking us to tilings of twist 00.
There is no fixed finite set of local moves connecting all domino tilings which works for any contractible cubiculated region ${\cal R}\subset{\mathbb{R}}^{3}$. It is conjectured that any two tilings of a box can be connected by a finite sequence of flips and trits. It is known for large classes of boxes that this holds for almost all tilings [6]. It is also known that any tiling of a box admits at least one flip or trit [2].
It is conjectured that for large boxes the twist follows a normal distribution. This is known to be true for boxes of size $L\times M\times N$ where $L,M\geq 4$ are even and fixed and $N$ tends to infinity [7]. In this case, it is also known that for almost all pairs $({\mathbf{t}}_{0},{\mathbf{t}}_{1})$ of tilings with $\operatorname{Tw}({\mathbf{t}}_{0})=\operatorname{Tw}({\mathbf{t}}_{1})$ there exists a finite sequence of flips connecting ${\mathbf{t}}_{0}$ and ${\mathbf{t}}_{1}$. This implies that connected components under flips almost coincide with level sets of the function $\operatorname{Tw}$. Equivalently, within a (large) level set of $\operatorname{Tw}$ there exists a giant component.
@@
The aim of the present paper is to investigate related questions for slab tilings of 3d regions. A slab is a $2\times 2\times 1$ cuboid. Let us draw a small example in Figure 2. As for domino tilings, we depict a slab tiling by floors, numbered from left to right. A slab which is contained in only one floor (horizontal) is represented as a $2\times 2$ light grey square. A slab which crosses two floors (vertical) is represented as two planar dominoes. These two dominoes must have the same projection on the first floor. The domino on the lower (left) floor is painted dark grey and the domino on the upper (right) floor is painted white.
Figure 2: A slab tiling of a $6\times 6\times 6$ box
@@
We define flips for slabs in the same way as for dominoes: we take two slabs which have a $2\times 2$ face in common, remove them from the region, and then place them back in one of the two other positions.
A natural question is to investigate connected components via flips of the set of slab tilings of a given region. The existence of the twist as an invariant under flips is crucial for the study of similar questions for domino tilings: does a similar invariant exist for slab tilings? As we shall see, the answer is yes, but there are many significant differences. The new invariant is the triple twist ($\operatorname{TTw}$), which assumes values in ${\mathbb{Z}}^{3}$. Its construction, which takes a large portion of the present paper, shows that it is related to the twist (for dominoes) and indeed invariant under flips. The triple twist assumes a large number of values.
Theorem 1.
Consider a cubiculated box ${\cal R}$ of size $N\times N\times N$, $N\in 2{\mathbb{N}}$. If $N$ is sufficiently large then there exists positive constants $C_{-}<C_{+}$ with the following properties. If ${\mathbf{t}}$ is a slab tiling of ${\cal R}$ then $|\operatorname{TTw}({\mathbf{t}})|<C_{+}N^{4}$. If $t\in(2{\mathbb{Z}})^{3}$ and $|t|<C_{-}N^{4}$ then there exists a tiling ${\mathbf{t}}$ of ${\cal R}$ with $\operatorname{TTw}({\mathbf{t}})=t$.
The previous theorem gives a lower bound for the number of connected components under flips of the set of slab tilings of a large box (see Proposition 23 for an explicit formula). We easily obtain an upper bound for the set of values of $\operatorname{TTw}$, but the number of connected components appears to be much larger (and harder to estimate). In some simple cases, such as a $4\times 4\times N$ or an $M\times N\times 2$ box, there is only one connected component.
The construction of the triple twist leads us to consider mixed tilings, consisting of horizontal slabs and vertical dominoes. Given a slab tiling, we construct a corresponding mixed tiling (Section 3). If the initial region is a box, this construction can be performed in any of the three canonical directions. Given in turn a mixed tiling, we use a four color pattern to obtain a corresponding domino tiling (Section 4): we then compute its twist. The triple twist is computed by performing the above construction in all three directions. As for the domino tilings, these new versions of the twist are well behaved under symmetries.
As for domino tilings, there exist examples of slab tilings which admit no flip (see Figures 20 and 21). There are similar examples of mixed tilings which admit no flip (see Figure 7). We find it more surprising to observe that there exists no fixed set of local moves which makes the set of slab tilings connected under such moves. More than that: given a fixed set of local moves (involving a bounded number of slabs) there exist slab tilings which admit none of these local moves (see Lemma 6.3, Corollary 6.4 and Remark 6.5).
In Section 2 we define and review basic properties of the twist for domino tilings. We introduce mixed tilings in Section 3. In Section 4 we consider a way of using four colors to paint the unit cubes in order to construct the twist invariants for mixed and slab tilings. In Section 5 we discuss some properties of these invariants, including the effect of symmetries. Section 6 includes examples and results concerning connectivity under local moves. Theorem 1 and related results are proved in Section 7. Finally, in Section 8 we list some open problems.
2 Construction of the twist
This section contains one among several known constructions of the twist for domino tilings of a cubiculated region ${\cal R}\subset{\mathbb{R}}^{3}$. This is included for completeness; we follow [4]. See [1, 3] for alternative constructions, sometimes more general.
@@ 2 Construction of the twist
The effect is a function which takes two dominoes in a domino tiling ${\mathbf{t}}$ and returns a value in $\{0,\pm 1\}$. On ${\mathbf{t}}$, we choose a horizontal domino $d_{0}$ such that the line through the centers of two cubes covered by $d_{0}$ is parallel to the base vector $\mathbf{e_{2}}$. Let $d_{1}$ be a vertical domino in ${\mathbf{t}}$. We say $d_{1}$ affects $d_{0}$ if any of the cubes of $d_{1}$ is on the same floor as $d_{0}$ and if the projections of the interiors of $d_{0}$ and $d_{1}$ over the plane $x=0$ have non-empty intersection. Figure 3 shows some examples for which there are effects. If these conditions do not hold we say the effect of $d_{1}$ on $d_{0}$ is 00.
@@
Figure 3: Examples of effects for pairs of dominoes
We then define the twist $\operatorname{Tw}({\mathbf{t}})$ as $\frac{1}{4}$ times the sum of all the effects of $d_{1}$ on $d_{0}$ for all the suitable pairs $(d_{0},d_{1})$. If ${\cal R}$ is a cylinder and ${\mathbf{t}}$ is a tiling of ${\cal R}$ then $\operatorname{Tw}({\mathbf{t}})$ is an integer [4].
For example, the twist of the tiling in Figure 1 is $-1$, since each floor has a contribution of $-\frac{1}{2}$. If we appy a trit to this tiling, the twist of the new tiling is 00. In Figure 4 we show two examples of tilings of cylinders, and in Figure 5 we rotate each of them, noting the twist is unchanged, consistently with the remarks in the Introduction.
Figure 4: Two examples of tilings of cylinders; in both cases, $\operatorname{Tw}=0$
Figure 5: The $x$-view of the two tilings in Figure 4, i.e., the same tilings after a rotation; we still have $\operatorname{Tw}=0$
3 Mixed tilings
The study of slab tilings led us to consider another class of tilings: mixed tilings. Given a three dimensional region ${\cal R}\subset{\mathbb{R}}^{3}$ with a specified vertical direction, a mixed tiling is a decomposition of ${\cal R}$ into horizontal slabs and vertical dominoes. We admit as mixed tilings the configurations where there are only horizontal slabs or only vertical dominoes.
Figure 6: A non trivial example of mixed tiling
The mixed tiling in Figure 6 can be obtained from the slab tiling in Figure 2 by slicing each vertical slab in half along a plane perpendicular to the paper. This process of slicing vertical slabs in half to transform a slab tiling into a mixed tiling can always be done. However, it is not always possible to undo this process in a well defined manner. For example, in Figure 6 there are many ways to join the vertical dominoes which cross the first and second floor. Also, there are mixed tilings which do not come from any slab tiling, as in Figure 7.
Figure 7: A mixed tiling which does not correspond to a slab tiling, obtained from the tiling in Figure 6 by taking the central $4\times 4\times 6$ box
We can also define a flip in a mixed tiling. In one way, we take two adjacent slabs which lie on different floors, perform a slab flip and then slice the resulting vertical slabs in half. In the other way, we take four vertical dominoes which form a $2\times 2\times 2$ cube, join two pairs of adjacent dominoes and perform a slab flip to end with two horizontal slabs. The mixed tiling in Figure 7 admits no flip; See Figure 8 for an example of a flip.
Figure 8: We show one flip performed on the tiling of Figure 6
@@ 4 Coloring and twist
Figure 9: A slab-type coloring of a $6\times 6\times 6$ box, based on Example 4.1
@@
Figure 10: The top left corner of the first two floors of ${\cal R}$
@@
The second step of the process is to transform the region. We first consider cylinders of the form ${\cal R}={\cal D}\times[0,N]$ and the $z$-view. A pair of colors $(c_{0},c_{1})$ is good (for this view) if $c_{0}$ and $c_{1}$ are on a same diagonal on the first floor: the good pairs are therefore red-green and blue-yellow. The colors in a good pair are also called good; the other colors are bad. Informally, we keep all cubes (or squares) of one of the two good colors and delete all cubes of bad colors. We then inflate the remaining cubes so as to obtain cuboids which fill in the space left by the deleted cubes. Figure 11 shows an example: the initial region is as in Figure 9 and the good pair is red-green.
Figure 11: The region in Figure 9 is transformed by deleting the blue and yellow cubes and inflating the red and green ones
@@
We are now left with the task of transforming the tilings. We first consider slab tilings; let ${\mathbf{t}}$ be a tiling of ${\cal R}$. Each slab $P$ of ${\mathbf{t}}$ covers two preserved cubes and two deleted cubes. Let $S_{0}$ and $S_{1}$ be the two preserved cubes, which are transformed into $\tilde{S}_{0}$ and $\tilde{S}_{1}$, respectively. We say that $P$ is taken to the domino $\tilde{P}$, the union of $\tilde{S}_{0}$ and $\tilde{S}_{1}$. See Figures 2 and 12 for an example. In the second row of Figure 12, we show the tiling rotated clockwise by $45^{\circ}$; in this example, we have $\operatorname{Tw}=+2$.
Figure 12: The domino tiling obtained from the slab tiling in Figure 2 or the mixed tiling in Figure 6
@@
Figure 13: The figures corresponding to the $y$-view, with the blue-green good pair; $\operatorname{Tw}(\tilde{\mathbf{t}})=0$
@@
Figure 14: The figures corresponding to the $x$-view, with the green-yellow good pair; $\operatorname{Tw}(\tilde{\mathbf{t}})=0$
@@
Figure 15: Two possibilities of flipping two slabs on the top left corner of the first floor of Figure 2; under the transformation, the first possibility yields the same domino tiling as in Figure 12
Figure 16: The domino tiling from the second possibility in Figure 15 after the transformation
@@ 5 Properties of the Triple Twist
Figure 17: An example of a mixed tiling; the top left cube on the first floor is $[0,1]^{3}$, which is green
@@
Figure 19: We have, in order: one instance of case 1, two instances of case 2 and two instances of case 3
@@
The following proposition is similar, but for slab tilings. A symmetry of ${\mathbb{R}}^{3}$ which preserves unit cubes is a map of the form $q(w)=\mathbf{v_{0}}+\mathbf{Qw}$ where $\mathbf{v_{0}}\in{\mathbb{Z}}^{3}$ and $\mathbf{Q}$ is a $3\times 3$ permutation matrix with signs. Let $\operatorname{sign}(q)=\det(\mathbf{Q})\in\{\pm 1\}$.
@@
6 Connectivity under local moves
In this section we return to our main problem: connectivity under local moves of spaces of tilings. We first show that the space of slab tilings is connected via flips for a few special regions. We then show that there exists no finite fixed set of local moves which works for slab tilings of arbitrary boxes. Equivalently, there is no bound on the number of slabs which have to participate in each move.
We begin with two-floored regions.